The average cholesterol content of a 100-gram bar of a certain brand of butter is 215 mg with a standard deviation of 15 mg. If a sample of twenty-five 100-gram bars of butter is selected, find the probability that the sample mean is less than 220 mg. Assume that the variable is normally distributed.



1
Expert's answer
2022-03-23T17:56:58-0400

"P(\bar X<220)=P(<\frac{220-215}{\frac{15}{\sqrt{25}}})=P(Z<1.67)=0.9525."